OpenStudy (anonymous):

Solve: 6x - 2(x + 1) = 3x - 10 A. 8 B. -8 C. -12 D. 12

OpenStudy (anonymous):

@Jacob902

OpenStudy (anonymous):

6x-2x-2=3x-10 6x-3x-2x=-10+2 x=-8

OpenStudy (anonymous):

Answer B/

OpenStudy (anonymous):

thnx guys :)

OpenStudy (anonymous):

or nvm lol just to @MBilbeisi

OpenStudy (anonymous):

Hahaha :)